The value of Ka for hydrocyanic acid is 4.00×10-10.
What is the value of Kb, for its conjugate base, CN-?
1
Expert's answer
2017-04-25T01:53:08-0400
The Ka for an acid and Kb for a conjugate base are related to each other according to the equation
K_a*K_b=10^(-14).
So, the value of Kb for CN is
K_b (CN^- )=(10^(-14))/(K_a (HCN) )=(10^(-14))/(4*10^(-10) )=2.5*10^(-5).
Answer: Kb (CN ) = 2.5*10-5.